Given the equation of the circle (x – 9)² + y² = 484, the center of the circle is located at __________, and its radius has a length of __________ units.
Answer by Edwin McCravy(20086)

The general statement for all circles is:

Given the equation of the circle (x – h)² + (y-k)² = r², 
the center of the circle is located at (h,k), and its 
radius has a length of r units.

You should learn that.

Your equation is 

(x – 9)² + y² = 484

Since the square root of 484 is 22 and since y can
be written as (y - 0), your equation can be written as

(x – 9)² + (y - 0)² = 22²

Now if we compare that with

(x – h)² + (y - k)² = r²

we see that h = 9, k = 0, and r = 22.

So we make those substitutions in the general statement 
about all circles, we get

Given the equation of the circle (x – 9)² + (y - 0)² = 22², 
the center of the circle is located at (9,0), and its 
radius has a length of 22 units. 

Edwin